An iron block of mass 2kg Fall from height 10m. After colliding with ground it looses 25% of energy.find temperature rise in block.(specific heat of iron =470j/kg)

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Anand
 
(1/4)mgh = m*s* delta T
Now
delta T = gh/4s
= 10* 10/470*4
Now solve this and you will get the temp rise.
 
Regards
Arun (askIITians forum expert)